7/3 over y = 1/3 over 6???
Answer by stanbon(75887)

7/3 over y = 1/3 over 6
----------------------------
(7/3)/y = (1/3)/6
----
(7/3y) = (1/18)
--------------
Cross-multiply
3y = 7*18
y = 7*6
y = 42
-------------
Cheers,
Stan H.